A 3.2 Hz continuous wave travels on a Slinky. If the wavelength is 0.49 m, what is the speed of waves on the Slinky (in m/s)?
1
Expert's answer
2019-04-23T11:13:04-0400
"T = \frac{1}{f} = \frac{1}{3.2 Hz} =\frac{1}{3.2} s"

"wave speed = \frac{distance}{time} = \frac{0.49 m}{\frac{1}{3.2}s} = 1.568 \frac{m}{s} = 1.6 \frac{m}{s}"